OpenStudy (anonymous):

the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30

OpenStudy (amistre64):

how will you know when they are the correct ones?

OpenStudy (anonymous):

idk thats y im asking

OpenStudy (amistre64):

\[ax+by=c\] \[\frac{a}{c}x+\frac bc y=\frac cc\] \[\frac{x}{c/a}+\frac{y}{c/b}=1\] then we can read the ints froom under neath

OpenStudy (amistre64):

so, xint = 30/15 yint = 30/-5

OpenStudy (amistre64):

theres other ways of course, but I find this one is neglicated .... so i try to give it some breathing room